How do you factor out x^6 - y^6 using the difference of two cubes. I solved it using using difference of two squares, but I can't work out using cubes.
Answer by lwsshak3(11628)

How do you factor out x^6 - y^6 using the difference of two cubes. I solved it using using difference of two squares, but I can't work out using cubes.
***
x^6-y^6=(x^2)^3-(y^2)^3
=(x^2-y^2)(x^4+x^2y^2+y^4)
=(x+y)(x-y)(x^4+x^2y^2+y^4)